darkknight:

how would I get this one into exponential standard form: 2/(1.5^((x/3)-3))

darkknight:

i got to 2 ((1.5^3)/(1.5^(x/3)))

darkknight:

I don't know if I did this correct

justjm:

You are on the right track but you could simplify it more, so you can get it to the form ab^x remember that 1.5^(x/3) can be written as 1.5^(1/3*x), which can further be written as ((1.5)^(1/3))^x. Then you can simplify further

darkknight:

ty

darkknight:

what do you do after that?

justjm:

So you said you got 2 times (1.5^3)/((1.5^(1/3))^x) \(\frac{2\left(1.5^{3}\right)}{\left(1.5^{\frac{1}{3}}\right)^{x}}\) you can simplify the numerator using your calculator to 6.75, and then move the bottom denominator to the numerator by taking the negative exponent \(6.75\left(1.5^{\frac{1}{3}}\right)^{-x}\) You can use your exponent rules to make it \(6.75\left(1.5^{-\frac{1}{3}}\right)^{x}\) \(6.75\left(\frac{1}{1.5^{\frac{1}{3}}}\right)^{x}\) Use your calculator to evaluate the \(\frac{1}{1.5^{\frac{1}{3}}}\) to a basic decimal then you should get something in the form of ab^x

darkknight:

thanks

darkknight:

I understand it

justjm:

That's good bro, keep it up. You're welcome (: